Jack invested $8000, part at 10% and part at 15%. If the total interest at the end of the year is $900, how much did he invest at 10%?
a)$7000
b)$5000
c)$6000
d)$4000

Answer by checkley71(8403)

.1x+.15(8000-x)=900
.1x+1200-.15x=900
-.05x=900-1200
-.05x=-300
x=-300/-.05
x=6000 invested @ 10% answer c)